OpenStudy (anonymous):

find the area of a rectangle 2 yards by 15/4

OpenStudy (johnweldon1993):

Area of rectangle = length x width...so \[\large 2 \times \frac{15}{4}\]

OpenStudy (johnweldon1993):

which would then be \[\large \frac{2 \times 15}{4} = \frac{30}{4} = 7.5 \text{yards}\] Hope that makes sense :)